OpenStudy (anonymous):

What is the value of y so that the ordered pair (2, y) is a solution to the equation2x+y=-2 ? A. –6 B. –2 C. D. 0

Parth (parthkohli):

All right, so if \((2,y)\) is an ordered pair of the equation, then we have \(x=2\) and we're looking to find the value of \(y\).

Parth (parthkohli):

Do you understand what I said?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

a

Parth (parthkohli):

Thus, we can check the value of \(y\) when \(x = 2\).

OpenStudy (anonymous):

Yes i understood

OpenStudy (anonymous):

-6?

Parth (parthkohli):

When \(x=2\), then the equation \(2x + y = -2\) simplifies to \(2(2) + y = -2 \Rightarrow 4 + y = -2\).

Parth (parthkohli):

Righty.

OpenStudy (anonymous):

Yes thank you for all of you're help its just one month untill finals so this will help me alot thank you so much :-)

Parth (parthkohli):

No problem.